OpenStudy (anonymous):

Need help seeing where i am wrong in this problem.. have tried it three times and its all wrong. thanks.

OpenStudy (anonymous):

Find the constant difference for a hyperbola with foci at F1 (0, -10) and F2 (0, 10) and the point on the hyperbola (6, 7.5). The constant difference is: _____.

OpenStudy (anonymous):

I used the distance formula.. \[\sqrt{(0-6)^2 + (-10-7.5)^2} - \sqrt{(0-6)^2 + (10-7.5)^2}\]

OpenStudy (anonymous):

\[\sqrt{36 + (-17.5)} - \sqrt{36 + 6.25}\]

OpenStudy (anonymous):

Please help.. im confused on where i am wrong.

OpenStudy (aravindg):

10-7.5=?

OpenStudy (anonymous):

It adds up to be.. \[\sqrt{18.5} - \sqrt{422.25}\]

OpenStudy (aravindg):

10-7.5=? :)

OpenStudy (anonymous):

Its 2.5

OpenStudy (aravindg):

Oh wait sorry. Let me check again.

OpenStudy (aravindg):

Why didnt u square -17.5?

OpenStudy (anonymous):

I did.. didnt i. I first added 36 + -17.5 then get sqrt 18.5

OpenStudy (anonymous):

then i did 36 + 6.25 for the next part and get sqrt 42.25

OpenStudy (anonymous):

then i would square both of them.. the first one doest square easliy but the second part does and that equals 6.5

OpenStudy (anonymous):

Am i solving this wrong?

OpenStudy (aravindg):

Yes.

OpenStudy (aravindg):

\[\sqrt{6^2+(-17..5)^2}\]

OpenStudy (anonymous):

Where? trying to see but dont see where.

OpenStudy (aravindg):

look ^

OpenStudy (aravindg):

Also 36+6.5=42.25

OpenStudy (anonymous):

But why would i square them again if i already i did. here sqrt (0−6)^2+(−10−7.5) ^2 and that got me 0-6 = -6^2 = 36 and -10 -7.5 = -17.5^2 = 306.25... wait is this part where i am wrong?

OpenStudy (anonymous):

the answer is 12 right? :)

OpenStudy (aravindg):

Yes it is :)

OpenStudy (anonymous):

thanks!! :)

OpenStudy (aravindg):

yw :)